OpenStudy (anonymous):

what are the intrcepts of the equation: 2x+(3/2)y+3z=6

OpenStudy (anonymous):

Kush

OpenStudy (anonymous):

2x+(3/2)y+3z=6 divide by 6,we get 2x/6+3y/12+3z/6=1 x/3+y/4+z/2=1 hence x-intercept is=3 y-intercept is=4 z -intercept is=2